# Minimal Degrees, Volume Growth, and Curvature Decay on Complete Kähler Manifolds

### v1

@@ 2. Preliminary results
2.1. Structure of Kähler manifolds with maximal volume growth and the tangent cones
We first recall the main result in [16], and prove some relevant facts that will be used later.
Theorem 2.1.
Let $(M^{n},g,p)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth. Let $\operatorname{AVR}(M,g)=\nu>0$. Then
@@ 2.1. Structure of Kähler manifolds with maximal volume growth and the tangent cones
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,1)}f_{i}\overline{f_{j}}=\delta_{ij},\quad f_{i}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$ (2.1)
@@
The proof of these conclusions is implicitly contained in [11, 16]. For reader’s convenience, we give a thorough explanation. (1) and (2) are contained in Section 4 in [16].
The properness is also contained in [11]. It was shown that there exists a constant $D=D(n,\nu)$ such that let $(g_{1},\cdots,g_{k})$ be a linearly independent basis of $\mathcal{O}_{P}(M)$, where $k=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{P}(M)-1$, such that
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,1)}g_{i}\overline{g_{j}}=\delta_{ij},\quad g_{i}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,k\}.$
@@
Take an $n$-tuple algebraically independent holomorphic functions with minimal degrees in $(g_{1},\cdots,g_{k})$, denoted by $(f_{1},\cdots,f_{n})$. As in section 4 in [16], $(f_{1},\cdots,f_{n})$ serves as a biholomorphism of $M$ onto $\mathbb{C}^{n}$, and any polynomial growth function is a polynomial of them.
Since any $g_{i}$ is a polynomial of these $(f_{1},\cdots,f_{n})$ with degree no larger than $D$ by the three-circle theorem, by a contradiction argument we get
$\min\limits_{\partial B(p,r)}\sum\limits_{i=1}^{n}|f_{i}|^{2}\geq c(n,\nu)r^{\frac{2}{D}}$ (2.2)
@@
For (3), given any two such global coordinates $(f_{1},\cdots,f_{n})$ and $(h_{1},\cdots,h_{n})$. Assume their degrees are arranged in the nondecreasing order, then the matrix $\left(\frac{\partial h_{i}}{\partial f_{j}}\right)$ and $\left(\frac{\partial f_{i}}{\partial h_{j}}\right)$ are both polynomial matrices of $(f_{1},\cdots,f_{n})$ and are inverse of each other. Therefore they must both be constant matrices. Then due to the normalization condition (2.1), these two matrices are both orthogonal. ∎
In fact, the minimal degrees of all coordinates in Theorem 2.1 (3) described above are just the refined minimal order $\overrightarrow{d_{\operatorname{min}}}$ defined in the introduction, and it has been shown in [16].
Corollary 2.2.
Conjecture 1.5 (1) holds.
@@
If $M$ is of nonmaximal volume growth. By Theorem 2 in [12], $\mathcal{O}_{P}(M)=\mathbb{C}$, i.e. $\overrightarrow{d_{\operatorname{min}}}=\overrightarrow{+\infty}$, which is independent of the choice of $p$.
If $M$ is of maximal volume growth, due to [16], there exists a strictly increasing sequence $1\leq d_{1}<d_{2}<\cdots$, so that $\mathcal{O}_{P}(M)=\mathbb{C}\oplus\left(\mathcal{O}_{d_{1}}(M)/\mathbb{C}\right)\oplus\left(\mathcal{O}_{d_{2}}(M)/\mathcal{O}_{d_{1}}(M)\right)\oplus\cdots$ as a complex vector space. For any $k\in\mathbb{N}$, pick a maximal linearly independent vectors $f_{k1},\cdots,f_{km_{k}}$ of $\mathcal{O}_{d_{k}}(M)$ so that they form a basis of $\mathcal{O}_{d_{k}}(M)/\mathcal{O}_{d_{k-1}}(M)$ as quotient of vector spaces and no element in the span is given by polynomials of $\mathcal{O}_{d_{k-1}}(M)$. Then the first $n$ functions $\{f_{11},\cdots,f_{1m_{1}},f_{21},\cdots\}$ form a global coordinate in Theorem 2.1 after normalization as (2.1).
@@
Remark 2.3.
(1) As can be seen in the proof, if $M$ is of maximal volume growth, Corollary 2.2 still holds without the nonsplitting condition on the universal cover.
(2) For simplicity, if $M$ is of maximal volume growth we just call the coordinate as in (2) in Theorem 2.1 a canonical coordinate on $M$, with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$. And we arrange them in non-decreasing order of their degrees.
Let $M_{\infty}$ be a tangent cone of infinity of $M$, whose structure has been studied in [14, 15, 9, 19, 18]. Let’s summarize these results in the following:
Theorem 2.4.
@@
(1) $(M_{\infty},p_{\infty},d_{\infty})$ is a complex manifold which is biholomorphic to $\mathbb{C}^{n}$. Moreover, it’s a metric Kähler cone with $BK\geq 0$” (see Section 5 in [19] for relevant definitions).
(2) There exists a complete Kähler-Ricci flow solution $h(t)$ on $M_{\infty}$ for $t\in[0,+\infty)$ with nonnegative bisectional curvature such that $(M_{\infty},p_{\infty},d_{\infty})$ is also the Gromov-Hausdorff limit of $h(t)$ as $t\rightarrow 0$. Moreover, $(M_{\infty},h(t))$ is a gradient Kähler-Ricci soliton with nonnegative bisectional curvature.
(3) $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{\infty})=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ for any $d\geq 0$. And there exists a n-tuple of polynomial growth functions $(z_{1\infty},\cdots,z_{n\infty})$ serving as a proper biholomorphism onto $\mathbb{C}^{n}$. Each $z_{i\infty}$ is of homogeneous degree $d_{\operatorname{min}}^{(i)}$, i.e. there exists a eigenfunction $\phi_{i}$ of the Laplacian operator on the link such that $z_{i\infty}=r^{d_{\operatorname{min}}^{(i)}}\phi_{i}$ on $M_{\infty}$. They satisfy
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p_{\infty},1)}z_{i\infty}\overline{z_{j\infty}}=\delta_{ij},\quad z_{i\infty}(p_{\infty})=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$ (2.3)
@@
$\mathchoice{{\vbox{\hbox{$\displaystyle\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\displaystyle\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\displaystyle\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p_{k},1)}z_{ik}\overline{z_{jk}}=\delta_{ij},\quad z_{ik}(p_{k})=0\quad\text{ for any }i,j\in\{1,\cdots,n\},k\in\mathbb{N}.$ (2.4)
@@
(6) The volume of the unit ball in $M_{\infty}$ is $\frac{\omega_{2n}}{d_{\operatorname{min}}^{(1)}\cdots d_{\operatorname{min}}^{(n)}}$. In particular, Conjecture 1.5 (2) holds.
@@
(1) and (2) were proved in Proposition 6.1 in [19].
For (3), for any $d\geq 0$, $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{\infty})=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ by Proposition 4.1 in [16]. Now we show the existence of such coordinate.
Firstly we claim that the homothetic vector field $r\frac{\partial}{\partial r}$ is holomorphic on $M_{\infty}$. By Theorem 1.9 there exists a long-time Kähler-Ricci flow solution $g(t)$ on $M$ such that $(M_{\infty},J_{\infty},p_{\infty},h(t))$ is the pointed Cheeger-Hamilton limit of a blow-down sequence of $g(t)$, i.e. $(M,J,p,g_{i}(t)=\frac{1}{t_{i}}g(t_{i}t)$ for some sequence $t_{i}\rightarrow+\infty$. In other words, $J$ converges to $J_{\infty}$ through a sequence of diffeomorphisms exhausted on $M_{\infty}$. By Claim 4.3 in [16], the complexification of $r\frac{\partial}{\partial r}$ is in the span of several holomorphic vector fields on $M_{\infty}$. Moreover, $(r\frac{\partial}{\partial r})\mathcal{O}_{d}(M_{\infty})\subset\mathcal{O}_{d}(M_{\infty})$ holds for any $d\geq 0$. The claim is confirmed.
Therefore, $r\frac{\partial}{\partial r}$ is a contracted holomorphic field on $M_{\infty}$, from a result in [25], $M_{\infty}$ is biholomorphic to $\mathbb{C}^{n}$. Now follow Section 4 in [16], replace the holomorphic vector field $X$ in the proof by $r\frac{\partial}{\partial r}$, we conclude the proof of (3) except the homogenity of these coordinates. This is also obvious since we can just choose the highest order terms of each coordinate functions and they certainly form a new biholomorphism onto $\mathbb{C}^{n}$.
The proof of (4) is the same as the proof of (3) in Theorem 2.1.
For (5), the mean value inequality [10] implies that,
@@
holds for any $1\leq i\leq n$ and $k\in\mathbb{N}$. By the three-circle theorem in [13] and Cheng-Yau gradient estimate [5],
@@
on $B_{p_{i}}(r)$ for any $1\leq i\leq n$, $k\in\mathbb{N}$ and $r\geq\frac{1}{2}$. Arzela-Ascoli theorem implies that $z_{ik}$ converges to a holomorphic function $z_{i\infty}$ of homogeneous degree no more than $d_{\operatorname{min}}^{(i)}$ uniformly on compact subsets. By (2.2) and the Gromov-Hausdorff convergence,
$\min\limits_{\partial B(p_{\infty},r)}\sum\limits_{i=1}^{n}|z_{i\infty}|^{2}\geq c(n,\nu)r^{\frac{2}{D(n,\nu)}}>0.$ (2.5)
@@
on $M_{\infty}$. Lift to $M_{k}$ for large $k$ and use the properness (2.4) of $(z_{1k},\cdots,z_{nk})$, it implies $|P(z_{1k},\cdots,z_{nk})|$ can be arbitrarily small on arbitrarily large ball in $\mathbb{C}^{n}$. Thus $P\equiv 0$. And $z_{i\infty}$ must of homogeneous degree $d_{\text{min}}^{(i)}$ by the dimension estimate in (3), because on $M_{\infty}$ the minimal degree of a n-tuple algebraically independent holomorphic homogeneous functions has to be $\overrightarrow{d_{\operatorname{min}}}$. Therefore, (4) implies that $(z_{1\infty},\cdots,z_{n\infty})$ must form a global coordinate as in (3).
For (6), the first statement was proved in the proof of Theorem 3 in page 13 in [18]. This concludes the proof of Conjecture 1.5 (2) in the maximal volume growth case. The case of nonmaximal volume growth is trivial. ∎
Remark 2.5.
(1) As can be seen in the proof, if $M$ is of maximal volume growth, Conjecture 1.5 (2) still holds without the nonsplitting condition on the universal cover.
(2) For simplicity, we just call the coordinate as in (2) in Theorem 2.4 a canonical coordinate on $M_{\infty}$, with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$. And we arrange them in non-decreasing order of their degrees.
@@
Suppose $(M_{i},p_{i})\rightarrow(M_{\infty},p_{\infty})$ in the sense of Gromov-Hausdorff convergence, as in Proposition 5 in [15], each $f_{ik}\rightarrow f_{i\infty}$ with $f_{i\infty}$ is of homoegeneous degree $d_{\operatorname{min}}^{(i)}$.
@@
Compare these we complete the proof. ∎
Recently, Chu-Hao [7] obtain an optimal rigidity of the dimension estimate for polynomial growth holomorphic functions. Now we give an alternative proof, which seems much simpler.
Theorem 2.7.
@@
From Corollary 2.6, $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)=\#\{(m_{1},\cdots,m_{n})\in\mathbb{Z}^{n}_{\geq 0}\mid\sum\limits_{i=1}^{n}m_{i}d^{(i)}_{\min}\leq d\}$. We know that $M$ is biholomorphically isometric to $\mathbb{C}^{n}$ if and only if $d^{(i)}_{\min}=1$ for all $i$. Suppose $M$ is not isometric to $\mathbb{C}^{n}$. Then $d^{(n)}_{\min}>1$, then
$\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ $=\#\{(m_{1},\cdots,m_{n})\in\mathbb{Z}^{n}_{\geq 0}\mid\sum\limits_{i=1}^{n}m_{i}d^{(i)}_{\min}\leq d\}$
$\leq\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\#\{(m_{1},\cdots,m_{n-1})\in\mathbb{Z}^{n-1}_{\geq 0}\mid\sum\limits_{i=1}^{n}m_{i}<d\}$
@@
with equality holds if and only $d^{(i)}_{\operatorname{min}}=1$ for $i\leq n-1$ and $1<d^{(n)}_{\operatorname{min}}<\frac{d}{d-1}$. This contradicts with our assumption.
Now suppose the universal cover $\widetilde{M}$ splits. By the proof of Corollary 7.1 in [11], $\widetilde{M}=M_{1}^{m_{1}}\times M_{2}^{n-m_{1}}$, where $M_{1},M_{2}$ are both simply connected and has nonnegative bisectional curvature. Also $M_{1}$ is of maximal volume growth. $\mathcal{O}_{d}(M)$ can be seen as the $G$-invariant subring of $\mathcal{O}_{d}(M_{1})$, $G$ is a compact Lie group acting on $M_{1}$.

### v2

@@ 2. Preliminary results
2.1. Structure of Kähler manifolds with maximal volume growth and tangent cones
We first collect several results about the strucutures of complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth and their tangent cones, and prove some relevant facts that will be used later.
The following is the main theorem in [Liu19].
Theorem 4.
Let $(M^{n},g)$ be a complete noncompact Kähler manifold with nonnegative bisectional curvature and maximal volume growth, $p\in M$. Let $\operatorname{AVR}(M,g)=\nu>0$. Then
@@ 2.1. Structure of Kähler manifolds with maximal volume growth and tangent cones
(2.6) $\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,1)}f_{i}\overline{f_{j}}=\delta_{ij},\quad f_{i}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$
@@
The proof of these conclusions is implicitly contained in [Liu16a, Liu19]. For reader’s convenience, we give a thorough explanation. The proof of (1) and (2) is contained in Section 4 in [Liu19].
The properness is also contained in [Liu16a]. It’s shown that there exists a constant $D=D(n,\nu)$ such that let $(g_{1},\cdots,g_{k})$ be a linearly independent basis of $\mathcal{O}_{D}(M)$, where $k=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{D}(M)-1$, such that
$\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p,1)}g_{i}\overline{g_{j}}=\delta_{ij},\quad g_{i}(p)=0\quad\text{ for any }i,j\in\{1,\cdots,k\}.$
@@
Take an $n$-tuple algebraically independent holomorphic functions with minimal degrees in $(g_{1},\cdots,g_{k})$, denoted by $(f_{1},\cdots,f_{n})$. Following the proof in section 4 in [Liu19], $(f_{1},\cdots,f_{n})$ serves as a biholomorphism of $M$ onto $\mathbb{C}^{n}$, and any polynomial growth function is a polynomial of them.
Since any $g_{i}$ is a polynomial of these $(f_{1},\cdots,f_{n})$ with degree at most $D$ by the three-circle theorem, by a contradiction argument we get
(2.7) $\min\limits_{\partial B(p,r)}\sum\limits_{i=1}^{n}|f_{i}|^{2}\geq c(n,\nu)r^{\frac{2}{D}}$
@@
For (3), given any two such global coordinates $(f_{1},\cdots,f_{n})$ and $(h_{1},\cdots,h_{n})$. Assume their degrees are both arranged in the nondecreasing order, then the matrix $\left(\frac{\partial h_{i}}{\partial f_{j}}\right)$ and $\left(\frac{\partial f_{i}}{\partial h_{j}}\right)$ are both polynomial matrices of $(f_{1},\cdots,f_{n})$ and are inverse of each other. Therefore they must both be constant matrices. Then due to the normalization condition (2.6), these two matrices are both orthogonal. ∎
In fact, the minimal degrees of the coordinate in Theorem 4 (3) described above are just the refined minimal order $\overrightarrow{d_{\operatorname{min}}}$ defined in the introduction, and it has been shown in [Liu19].
Corollary 2.
Conjecture 2 (1) holds.
@@
If $M$ is of nonmaximal volume growth. By Theorem 2 in [Liu16b], $\mathcal{O}_{P}(M)=\mathbb{C}$, i.e. $\overrightarrow{d_{\operatorname{min}}}=\overrightarrow{+\infty}$, which is independent of the choice of $p$.
If $M$ is of maximal volume growth, due to [Liu19], there exists a strictly increasing sequence $1\leq d_{1}<d_{2}<\cdots$, so that $\mathcal{O}_{P}(M)=\mathbb{C}\oplus\left(\mathcal{O}_{d_{1}}(M)/\mathbb{C}\right)\oplus\left(\mathcal{O}_{d_{2}}(M)/\mathcal{O}_{d_{1}}(M)\right)\oplus\cdots$ as a complex vector space. For any $k\in\mathbb{N}$, pick a maximal linearly independent vectors $f_{k1},\cdots,f_{km_{k}}$ of $\mathcal{O}_{d_{k}}(M)$ so that they form a basis of $\mathcal{O}_{d_{k}}(M)/\mathcal{O}_{d_{k-1}}(M)$ as quotient of vector spaces and no element in the span is given by polynomials of $\mathcal{O}_{d_{k-1}}(M)$. Then the first $n$ functions $\{f_{11},\cdots,f_{1m_{1}},f_{21},\cdots\}$ form a global coordinate in Theorem 4 after normalization as (2.6).
@@
Remark.
(1) As can be seen in the proof, if $M$ is of maximal volume growth, Corollary 2 still holds without the nonsplitting condition on the universal cover.
(2) For simplicity, if $M$ is of maximal volume growth we just call the coordinate as in (2) in Theorem 4 a canonical coordinate on $M$, with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$. And we arrange them in non-decreasing order of their degrees.
Now we recall the definition of $\operatorname{BK}\geq K$” in Lott [Lott21].
Definition 2.
Let $X$ be a reduced complex space, then for each $x\in X$, there is a neighborhood $U_{x}$ of $x$ and an embedding $e_{x}:U_{x}\rightarrow\mathbb{C}^{N_{x}}$ so that $e\left(U_{x}\right)$ is the zero set of a finite number of analytic functions defined on an open set $V_{x}\subset\mathbb{C}^{N_{x}}$.
(1) We call $X$ a Kähler space if on each $U_{x}$, there exists a continuous plurisubharmonic function $\phi_{x}$, so that $\phi_{x}-\phi_{y}$ is pluriharmonic on each $U_{x}\cap U_{y}\neq\varnothing$. Thus $\omega=\sqrt{-1}\partial\bar{\partial}\phi_{x}$ on $U_{x}$ is a well-defined positive (1,1)-current, and we call it the Kähler current on $X$.
(2) Let $d$ be a metric on a Kähler space $(X,\omega)$, we call $(X,\omega,d)$ a metric Kähler space, if the metric $d$ induces the topology of $X$, so that if $\Sigma$ is an embedded holomorphic disk, then for all $x\in X$, $\omega|_{\Sigma}$ equals the 2-dimensional Hausdorff measure on each $\Sigma\cap U_{x}\neq\varnothing$.
(3) Given $(X,\omega,d)$ a metric Kähler space and $K\in\mathbb{R}$, let $d_{p}\in C(X)$ be the distance function from $p\in X$. Define $d_{p,K}\geq 0$ by
$d_{p,K}^{2}=\begin{cases}-\frac{4}{K}\log\cos\left(d_{p}\sqrt{\frac{K}{2}}\right)&\text{ if }K>0,\\ d_{p}^{2}&\text{ if }K=0,\\ \frac{4}{-K}\log\cosh\left(d_{p}\sqrt{\frac{-K}{2}}\right)&\text{ if }K<0.\end{cases}$
(If $K>0$, then we restrict to $d_{p}\leq\frac{\pi}{\sqrt{2K}}$.)
Then we call $X$ has $\operatorname{BK}\geq K$” if for any $x\in X$, $\phi_{x}-d_{p,K}^{2}/2$ is plurisubharmonic on $U_{x}$ for any $p\in X$.
Moreover, if $X$ is a smooth Kähler manifold, the condition $\operatorname{BK}\geq K$” is equivalent to that bisectional curvature has lower bound $K$.
The structure of tangent cones of infinity of $M$ has been studied in [Liu18a, Liu18b, LT21, Lott21, Liu24]. Let’s summarize these results in the following:
Theorem 5.
@@
(1) $(M_{\infty},p_{\infty},d_{\infty})$ is a complex manifold which is biholomorphic to $\mathbb{C}^{n}$. Moreover, it has $BK\geq 0$” with Kähler current $\omega=\frac{\sqrt{-1}}{2}\partial\bar{\partial}r^{2}$, where $r$ is the radial function on $M_{\infty}$.
(2) There exists a complete family of gradient expanding Kähler-Ricci soliton coming out of $M_{\infty}$, i.e. there exists $(M_{\infty},h(t))$ a family of gradient expanding Kähler-Ricci soliton with nonnegative bisectional curvature such that $h(t)\rightarrow d_{\infty}$ as $t\rightarrow 0$ in the Gromov-Hausdorff sense.
(3) $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{\infty})=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ for any $d\geq 0$. And there exists a n-tuple of polynomial growth functions $(z_{1\infty},\cdots,z_{n\infty})$ serving as a proper biholomorphism onto $\mathbb{C}^{n}$. Each $z_{i\infty}$ is of homogeneous degree $d_{\operatorname{min}}^{(i)}$. They satisfy
(2.8) $\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p_{\infty},1)}z_{i\infty}\overline{z_{j\infty}}=\delta_{ij},\quad z_{i\infty}(p_{\infty})=0\quad\text{ for any }i,j\in\{1,\cdots,n\}.$
@@
(2.9) $\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{B(p_{k},1)}z_{ik}\overline{z_{jk}}=\delta_{ij},\quad z_{ik}(p_{k})=0\quad\text{ for any }i,j\in\{1,\cdots,n\},k\in\mathbb{N}.$
@@
(6) The volume of the unit ball in $M_{\infty}$ is $\frac{\omega_{2n}}{d_{\operatorname{min}}^{(1)}\cdots d_{\operatorname{min}}^{(n)}}$. In particular, Conjecture 2 (2) holds.
@@
(1) and (2) are proved in Proposition 6.1 in [Lott21].
For (3), for any $d\geq 0$, $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M_{\infty})=\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ by Proposition 4.1 in [Liu19]. Next we show the existence of such coordinate.
Firstly we claim that the homothetic vector field $r\frac{\partial}{\partial r}$ is holomorphic on $M_{\infty}$. By Theorem 2 (1), (5) there exists a long-time Kähler-Ricci flow solution $g(t)$ on $M$ such that $(M_{\infty},J_{\infty},p_{\infty},h(t))$ is the pointed Cheeger-Hamilton limit of a blow-down sequence of $g(t)$, i.e. $(M,J,p,g_{i}(t)=\frac{1}{t_{i}}g(t_{i}t))$ for some sequence $t_{i}\rightarrow+\infty$. In other words, $J$ converges to $J_{\infty}$ through a sequence of diffeomorphisms exhausted on $M_{\infty}$. By Claim 4.3 in [Liu19], the complexification of $r\frac{\partial}{\partial r}$ is in the span of several holomorphic vector fields on $M_{\infty}$. Moreover, $(r\frac{\partial}{\partial r})\mathcal{O}_{d}(M_{\infty})\subset\mathcal{O}_{d}(M_{\infty})$ holds for any $d\geq 0$. The claim is confirmed.
Therefore, $r\frac{\partial}{\partial r}$ is a contracted holomorphic field on $M_{\infty}$, from a result in [RR88], $M_{\infty}$ is biholomorphic to $\mathbb{C}^{n}$. Now follow Section 4 in [Liu19], replace the holomorphic vector field $X$ in the proof by $r\frac{\partial}{\partial r}$, we conclude the proof of (3) except the homogenity of these coordinate functions. This is also obvious since we can just choose the highest order terms of each coordinate functions and they certainly form a new biholomorphism onto $\mathbb{C}^{n}$.
The proof of (4) is the same as the proof of (3) in Theorem 4, we just omit it here.
For (5), the mean value inequality [LS84] implies that,
@@
holds for any $1\leq i\leq n$ and $k\in\mathbb{N}$. By the three-circle theorem in [Liu16c] and Cheng-Yau gradient estimate [CY75],
@@
on $B_{p_{i}}(r)$ for any $1\leq i\leq n$, $k\in\mathbb{N}$ and $r\geq\frac{1}{2}$. Arzela-Ascoli theorem implies that $z_{ik}$ converges to a holomorphic function $z_{i\infty}$ of homogeneous degree at most $d_{\operatorname{min}}^{(i)}$ uniformly on compact subsets. By (2.7) and the Gromov-Hausdorff convergence,
(2.10) $\min\limits_{\partial B(p_{\infty},r)}\sum\limits_{i=1}^{n}|z_{i\infty}|^{2}\geq c(n,\nu)r^{\frac{2}{D(n,\nu)}}>0.$
@@
on $M_{\infty}$. Lift to $M_{k}$ for large $k$ and use the properness (2.9) of $(z_{1k},\cdots,z_{nk})$, it implies $|P(z_{1k},\cdots,z_{nk})|$ can be arbitrarily small on arbitrarily large ball in $\mathbb{C}^{n}$. Thus $P\equiv 0$. And $z_{i\infty}$ must of homogeneous degree $d_{\text{min}}^{(i)}$ by the dimension estimate in (3), because on $M_{\infty}$ the minimal degree of a n-tuple algebraically independent holomorphic homogeneous functions has to be $\overrightarrow{d_{\operatorname{min}}}$. Therefore $(z_{1\infty},\cdots,z_{n\infty})$ must form a global coordinate as in (3).
For (6), the first statement is proved in the proof of Theorem 3 in page 13 in [Liu24]. This concludes the proof of Conjecture 2 (2) in the maximal volume growth case. The case of nonmaximal volume growth is trivial. ∎
Remark.
(1) As can be seen in the proof, if $M$ is of maximal volume growth, Conjecture 2 (2) still holds without the nonsplitting condition on the universal cover.
(2) For simplicity, we just call the coordinate as in (2) in Theorem 5 a canonical coordinate on $M_{\infty}$, with degrees $\left(d_{\operatorname{min}}^{(1)},\cdots,d_{\operatorname{min}}^{(n)}\right)$. And we arrange them in non-decreasing order of their degrees.
@@
Suppose $(M_{i},p_{i})\rightarrow(M_{\infty},p_{\infty})$ in the sense of Gromov-Hausdorff convergence, as in Proposition 5 in [Liu18b], each $f_{ik}\rightarrow f_{i\infty}$ with $f_{i\infty}$ is of homoegeneous degree $d_{\operatorname{min}}^{(i)}$.
@@
Comparing these we complete the proof. ∎
Recently, Chu-Hao [CH25] obtained an optimal rigidity of the dimension estimate for polynomial growth holomorphic functions. Now we give an alternative proof, which seems much simpler.
Theorem 6 (Theorem 1.6 in [CH25]).
@@
From Corollary 3, $\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)=\#\{(m_{1},\cdots,m_{n})\in\mathbb{N}^{n}\mid\sum\limits_{i=1}^{n}m_{i}d^{(i)}_{\min}\leq d\}$. We know that $M$ is biholomorphically isometric to $\mathbb{C}^{n}$ if and only if $d^{(i)}_{\min}=1$ for all $i$. Suppose $M$ is not isometric to $\mathbb{C}^{n}$. Then $d^{(n)}_{\min}>1$, then
$\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(M)$ $=\#\{(m_{1},\cdots,m_{n})\in\mathbb{N}^{n}\mid\sum\limits_{i=1}^{n}m_{i}d^{(i)}_{\min}\leq d\}$
$\leq\operatorname{dim}_{\mathbb{C}}\mathcal{O}_{d}(\mathbb{C}^{n})-\#\{(m_{1},\cdots,m_{n-1})\in\mathbb{N}^{n-1}\mid\sum\limits_{i=1}^{n}m_{i}<d\}$
@@
with equality holds if and only if $d^{(i)}_{\operatorname{min}}=1$ for $i\leq n-1$ and $1<d^{(n)}_{\operatorname{min}}<\frac{d}{d-1}$. This contradicts with our assumption.
Now suppose the universal cover $\widetilde{M}$ splits. By the proof of Corollary 7.1 in [Liu16a], $\widetilde{M}=M_{1}^{m_{1}}\times M_{2}^{n-m_{1}}$, where $M_{1},M_{2}$ are both simply connected and has nonnegative bisectional curvature. Also $M_{1}$ is of maximal volume growth. $\mathcal{O}_{d}(M)$ can be seen as the $G$-invariant subring of $\mathcal{O}_{d}(M_{1})$, $G$ is a compact Lie group acting on $M_{1}$.